OpenStudy (anonymous):

What is the simplified form of the quantity y-squared minus y minus 12 over the quantity y-squared minus 2y minus 15?

OpenStudy (anonymous):

y^2 - y -12/y^2 - 2y - 15

OpenStudy (dan815):

factor

OpenStudy (anonymous):

\[\frac{ y^{2}-y-12 }{ y^{2}-2y-15 }\]

OpenStudy (anonymous):

factorize the numerator and the denominator

OpenStudy (anonymous):

\[\frac{ (y-4)*(y+3) }{ (y-5)*(y+3) }\]

OpenStudy (anonymous):

the (y+3) cancel out leaving you with \[\frac{ y-4 }{ y-5 }\]

OpenStudy (anonymous):

y cannot equal -3 or 5 though, or else the denominator is 0

OpenStudy (anonymous):

So when you factorized: y - 4 * y + 3 is equal to -12 y - 5 * y + 3 is equal to -15 which is shown in the original equation. hmmmm... ok thank you!! :)

OpenStudy (anonymous):

(y-4)*(y+3) does not EQUAL -12, it equals y^2 -y -12